Please demonstrate how to solve the following equations by the substitution method:
y=x+3
2x+y=15

Answer by Alan3354(69443)

Please demonstrate how to solve the following equations by the substitution method:
y=x+3
2x+y=15
---------
Since y is already isolated in the 1st eqn, sub for y into the 2nd eqn.
2x + y = 15
2x + (x+3) = 15
3x + 3 = 15
3x = 12
x = 4
-------
y = x+3
y = 7